Derive the equation of motion of a simple harmonic motion and obtain the expression for its velocity time period and frequency

1
Expert's answer
2022-03-25T09:11:03-0400

"\ddot x+\frac km x=0,"

"\omega^2={\frac km},"

"\ddot x+\omega^2 x=0,"

"\omega=\sqrt{\frac km},"

"T=\frac{2\pi}{\omega}=2\pi\sqrt{\frac mk}."